# Each of those eight coalition logics is also determined by six other kinds of models

### v1

@@ 3.1 Eight classes of grand-coalition-first action models
Definition 9 (Outcome functions and availability functions).
@@
Outcome functions in grand-coalition-first action models are given in a from-whole-to-part way: the outcome states of a coalition $\mathrm{C}$ performing a joint action $\sigma_{\mathrm{C}}$ at a state $s$ are determined by the outcome states of the grand coalition performing action profiles containing $\sigma_{\mathrm{C}}$ at $s$.
@@
Then, for all $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $\mathtt{av}_{\mathrm{C}}(s)=\{\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}\mid\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}\text{ for some }\sigma_{\mathtt{AG}}\in\mathtt{av}_{\mathtt{AG}}(s)\}$.
This fact indicates that availability functions in grand-coalition-first action models also are given in a from-whole-to-part way: available joint actions of a coalition $\mathrm{C}$ at a state are determined by available joint actions the grand coalition at the state.
Definition 10 (Successor functions).
Define the successor function $\mathtt{suc}$ as follows: for all $s\in\mathtt{ST}$,
For any set $X$ of states, we say $\Delta\subseteq\mathcal{P}(X)$ is a general cover of $X$ if $\bigcup\Delta=X$. A general cover $\Delta$ is a cover if $\emptyset\notin\Delta$.
Fact 3.
@@
Proof.
$\Rightarrow$
Let $\mathrm{C}\subseteq\mathtt{AG}$. Let $w\in\bigcup\{\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\mid\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}\}$. Then, there is $\sigma_{\mathrm{C}}$ such that $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$ and $w\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$. Then, there is $\sigma_{\mathtt{AG}}$ such that $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$, $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}$ and $w\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})$. Then, $w\in\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\mid\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}\}$. That is $w\in\mathtt{suc}(s)$.
$\Leftarrow$
Let $w\in\mathtt{suc}(s)$. Then, there is $\sigma_{\mathtt{AG}}$ such that $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$ and $w\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})$. Let $\sigma_{\mathtt{AG}}=\sigma_{\mathrm{C}}\cup\sigma_{\overline{\mathrm{C}}}$. Then, $w\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$. Then, $w\in\bigcup\{\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\mid\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}\}$.
@@
  • $\mathtt{G}\textbf{-}\mathtt{AM}$ is independent if for all $s\in\mathtt{ST}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}(s)$, and $\sigma_{{\mathrm{D}}}\in\mathtt{JA}_{\mathrm{D}}(s)$, if $\sigma_{{\mathrm{C}}}\in\mathtt{av}_{\mathrm{C}}(s)$, and $\sigma_{{\mathrm{D}}}\in\mathtt{av}_{\mathrm{D}}(s)$, then $\sigma_{{\mathrm{C}}}\cup\sigma_{{\mathrm{D}}}\in\mathtt{av}_{\mathrm{C}\cup\mathrm{D}}(s)$;
@@
We use $\mathrm{ES}$ to indicate the set of the eight strings: $\epsilon$, $\mathtt{S}$, $\mathtt{I}$, $\mathtt{D}$, $\mathtt{SI}$, $\mathtt{SD}$, $\mathtt{ID}$, and $\mathtt{SID}$. We let the three symbols $\mathtt{S}$, $\mathtt{I}$ and $\mathtt{D}$ correspond to the three properties, respectively. We use the eight strings $\epsilon$, $\mathtt{S}$, $\mathtt{I}$, $\mathtt{D}$, $\mathtt{SI}$, $\mathtt{SD}$, $\mathtt{ID}$, and $\mathtt{SID}$ to correspond to the eight combinations of the three properties, respectively.
For every $\mathtt{X}\in\mathrm{ES}$ and grand-coalition-first action model $\mathtt{G}\textbf{-}\mathtt{AM}$, we say $\mathtt{G}\textbf{-}\mathtt{AM}$ is a grand-coalition-first action $\mathtt{X}$-model if $\mathtt{G}\textbf{-}\mathtt{AM}$ has the properties corresponding to $\mathtt{X}$.
@@ 3.2 Eight axiomatic systems
Independence of agents ($\mathtt{A}\text{-}\mathtt{IA}$): $(\langle[\mathrm{C}]\rangle\phi\land\langle[\mathrm{D}]\rangle\psi)\rightarrow\langle[\mathrm{C}\cup\mathrm{D}]\rangle(\phi\land\psi)$
where $\mathrm{C}\cap\mathrm{D}=\emptyset$
@@
For all $\mathtt{X}$ in $\mathrm{ES}$, the axiomatic system for $\mathsf{MCL+X}$ consists of the axioms and inferences of $\mathsf{MCL}$, and the axioms corresponding to the elements of $\mathtt{X}$.
@@
The axiomatic system for $\mathsf{MCL+X}$ given in Definition 13 is sound and complete with respect to the set of valid formulas in $\Phi$ relative to the class of grand-coalition-first action $\mathtt{X}$-models.
@@
The class of grand-coalition-first $\mathsf{SID}$-models is $\alpha$-representable ([11, 5]), and the class of grand-coalition-first $\mathsf{SD}$-models is $\alpha$-representable ([14]). Are the other six classes of grand-coalition-first action models $\alpha$-representable? Are the eight classes of grand-coalition-first action models $z$-representable? It is not known yet.
4 About compositionality of availability functions and outcome functions in grand-coalition-first action models
We say availability functions of a grand-coalition-first action model is compositional if the availability function of a coalition can be derived from the availability functions of its members in a natural way. In a similar sense, we talk about compositionality of outcome functions of grand-coalition-first action models.
4.1 Neither outcome functions nor availability functions in grand-coa-lition-first action models are compositional
The following two facts are easy to verify, and we skip their proofs.
Fact 4.
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action model, $\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of availability functions of $\mathtt{G}\textbf{-}\mathtt{AM}$, $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{G}\textbf{-}\mathtt{AM}$, and $s\in\mathtt{ST}$.
Then:
  1. (1)
    1. (a)
    2. (b)
  2. (2)
    1. (a)
      for all $s\in\mathtt{ST}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}(s)$, and $\sigma_{{\mathrm{D}}}\in\mathtt{JA}_{\mathrm{D}}(s)$, if $\sigma_{{\mathrm{C}}}\cup\sigma_{{\mathrm{D}}}\in\mathtt{av}_{\mathrm{C}\cup\mathrm{D}}(s)$, then $\sigma_{{\mathrm{C}}}\in\mathtt{av}_{\mathrm{C}}(s)$ and $\sigma_{{\mathrm{D}}}\in\mathtt{av}_{\mathrm{D}}(s)$;
    2. (b)
      for all $s\in\mathtt{ST}$, $\mathrm{C}\subseteq\mathtt{AG}$, where $\mathrm{C}=\{a_{1},\dots,a_{n}\}$, and $\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}(s)$, where $\sigma_{\mathrm{C}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$, if $\sigma_{{\mathrm{C}}}\in\mathtt{av}_{\mathrm{C}}(s)$, then $\sigma_{{a_{1}}}\in\mathtt{av}_{a_{1}}(s),\dots,\sigma_{{a_{n}}}\in\mathtt{av}_{a_{n}}(s)$.
Note that none of the converses of these items holds. Figure 1 indicates a counter-example for the first two items and Figure 2 indicates a counter-example for the last two items.
$s_{0}$$s_{1}$$s_{2}$$\{p,q\}$$\{p\}$$\{q\}$$\alpha_{1}\beta_{1}\atop\alpha_{2}\beta_{2}$$\alpha_{1}\beta_{2}\atop\alpha_{2}\beta_{1}$$\alpha_{1}\beta_{1}$$\alpha_{1}\beta_{1}$
Figure 1: An example of that outcome functions in grand-coalition-first action models are not generally compositional. Here, we suppose $\mathtt{AG}=\{a,b\}$.
$s_{0}$$s_{1}$$s_{2}$$\{p,q\}$$\{p\}$$\{q\}$$\alpha_{1}\beta_{2}$$\alpha_{2}\beta_{1}$$\alpha_{1}\beta_{1}$$\alpha_{1}\beta_{1}$
Figure 2: An example of that availability functions in grand-coalition-first action models are not generally compositional. Here, we suppose $\mathtt{AG}=\{a,b\}$.
4.2 Availability functions in grand-coalition-first action models meeting independence of agents are compositional
The following fact is easy to check:
Fact 5.
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a grand-coalition-first action $\mathtt{I}$-model.
Then:
  1. (1)
    for all $s\in\mathtt{ST}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}(s)$ and $\sigma_{{\mathrm{D}}}\in\mathtt{JA}_{\mathrm{D}}(s)$: $\sigma_{{\mathrm{C}}}\cup\sigma_{{\mathrm{D}}}\in\mathtt{av}_{\mathrm{C}\cup\mathrm{D}}(s)$ if and only if $\sigma_{{\mathrm{C}}}\in\mathtt{av}_{\mathrm{C}}(s)$ and $\sigma_{{\mathrm{D}}}\in\mathtt{av}_{\mathrm{D}}(s)$;
  2. (2)
    for all $s\in\mathtt{ST}$, $\mathrm{C}\subseteq\mathtt{AG}$, where $\mathrm{C}=\{a_{1},\dots,a_{n}\}$, and $\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}(s)$, where $\sigma_{\mathrm{C}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$: $\sigma_{{\mathrm{C}}}\in\mathtt{av}_{\mathrm{C}}(s)$ if and only if $\sigma_{{a_{1}}}\in\mathtt{av}_{a_{1}}(s),\dots,\sigma_{{a_{n}}}\in\mathtt{av}_{a_{n}}(s)$.
Remarks
Models of Coalition Logic are grand-coalition-first action $\mathtt{SID}$-models. In those models, availability functions are compositional, but outcome functions might not be. Figure 1 indicates a counter-example.
@@ 4.3 Outcome functions are compositional in many situations
There are many situations where every agent can change a component of the state, and different components of the state that can be changed by different agents are independent of each other. Outcome functions are compositional in these situations.
@@
Adam and Bob control a ship lock, which has two doors. Adam can close or open the front door and Bob can close or open the back door. They cannot open the two doors at the same time. This scenario can be represented by the grand-coalition-first action model depicted in Figure 3.
$s_{1}$$s_{2}$$s_{3}$$\{c_{f},c_{b}\}$$\{c_{f}\}$$\{c_{b}\}$$(\mathtt{skip},\mathtt{skip})$$(\mathtt{skip},\mathtt{skip})$$(\mathtt{skip},\mathtt{skip})$$(\mathtt{skip},\mathtt{open_{b}})$$(\mathtt{skip},\mathtt{close_{b}})$$(\mathtt{open_{f}},\mathtt{skip})$$(\mathtt{close_{f}},\mathtt{skip})$
Figure 3: This figure indicates a grand-coalition-first action model for the situation of Example 1, where outcome functions are compositional. Here $c_{f}$ and $c_{b}$, respectively, express the propositions the front door is closed and the back door is closed. Here are some instances of outcome functions: (1) $\mathsf{out}_{\{a,b\}}(s_{1},(\mathtt{skip},\mathtt{skip}))=\mathsf{out}_{a}(s_{1},\mathtt{skip})\cap\mathsf{out}_{b}(s_{1},\mathtt{skip})=\{s_{1},s_{2}\}\cap\{s_{1},s_{3}\}=\{s_{1}\}$; (2) $\mathsf{out}_{\{a,b\}}(s_{2},(\mathtt{close_{f}},\mathtt{skip}))=\mathsf{out}_{a}(s_{2},\mathtt{close_{f}})\cap\mathsf{out}_{b}(s_{2},\mathtt{skip})=\{s_{2}\}\cap\{s_{1},s_{2}\}=\{s_{2}\}$; (3) $\mathsf{out}_{\{a,b\}}(s_{3},(\mathtt{close_{f}},\mathtt{skip}))=\mathsf{out}_{a}(s_{3},\mathtt{close_{f}})\cap\mathsf{out}_{b}(s_{3},\mathtt{skip})=\{s_{1}\}\cap\{s_{1},s_{3}\}=\{s_{1}\}$.
@@
A system has two processes, $a$ and $b$. The process $a$ can assign values $1$ or 00 to the variable $x$, and the process $a$ can assign values $1$ or 00 to the variable $y$. When $x=0$, then $a$ can leave the value of $x$ unchanged or change it to $1$. When $x=1$, then $a$ leaves the value of $x$ unchanged. Similarly, when $y=0$, then $b$ can leave the value of $y$ unchanged or change it to $1$. When $y=1$, then $b$ leaves the value of $y$ unchanged. Figure 4 indicates a grand-coalition-first model for this scenario, where outcome functions are compositional.
@@
Figure 4: This figure indicates a grand-coalition-first action model where outcome functions are compositional. Here are some instances: (1) $\mathsf{out}_{\{a,b\}}(s_{1},(\mathtt{skip},\mathtt{skip}))=\mathsf{out}_{a}(s_{1},\mathtt{skip})\cap\mathsf{out}_{b}(s_{1},\mathtt{skip})=\{s_{1}\}\cap\{s_{1}\}=\{s_{1}\}$; (2) $\mathsf{out}_{\{a,b\}}(s_{2},(\mathtt{skip},\mathtt{y:=1}))=\mathsf{out}_{a}(s_{1},\mathtt{skip})\cap\mathsf{out}_{b}(s_{1},\mathtt{y:=1})=\{s_{1},s_{2}\}\cap\{s_{1}\}=\{s_{1}\}$; (3) $\mathsf{out}_{\{a,b\}}(s_{1},(\mathtt{x:=1},\mathtt{y:=1}))=\mathsf{out}_{a}(s_{1},\mathtt{x:=1})\cap\mathsf{out}_{b}(s_{1},\mathtt{y:=1})=\{s_{1},s_{2}\}\cap\{s_{1},s_{3}\}=\{s_{1}\}$.
The following settings about agency can be found in the literature, such as [6] and [7]: there are some agents; every atomic proposition is controlled by at most one agent. In these settings, outcome functions are compositional.

### v2

@@ 3.1 Eight classes of grand-coalition-first action models
Definition 9 (Outcome functions and availability functions of grand-coalition-first action models).
@@
Definition 10 (Successor functions of grand-coalition-first action models).
Define the successor function $\mathtt{suc}$ of $\mathtt{G}\textbf{-}\mathtt{AM}$ as follows: for all $s\in\mathtt{ST}$,
For every set $X$ of states, we say $\Delta\subseteq\mathcal{P}(X)$ is a general cover of $X$ if $\bigcup\Delta=X$. A general cover $\Delta$ is a cover if $\emptyset\notin\Delta$. It is easy to verify the following fact:
@@
Fact 3.
Then:
  1. (1)
  2. (2)
Note that neither of the converses of the two statements holds. Figure 1 indicates a counterexample for them.
$s_{0}$$s_{1}$$s_{2}$$\{p,q\}$$\{p\}$$\{q\}$$(\alpha_{1},\beta_{1})\atop(\alpha_{2},\beta_{2})$$(\alpha_{1},\beta_{2})\atop(\alpha_{2},\beta_{1})$$(\alpha_{1},\beta_{1})$$(\alpha_{1},\beta_{1})$
Figure 1: This figure is used to show that in grand-coalition-first action models, the set of outcome states of a joint action might not contain the intersection of the sets of outcome states of the joint action’s parts. A grand-coalition-first action model is indicated in the figure, where we suppose $\mathtt{AG}=\{a,b\}$. It can be seen: (1) $\mathsf{out}_{a}(s_{0},\alpha_{1})$, the set of outcome states of $a$ performing $\alpha_{1}$ at $s_{0}$, equals to $\{s_{1},s_{2}\}$, (2) $\mathsf{out}_{a}(s_{0},\beta_{1})$, the set of outcome states of $b$ performing $\beta_{1}$ at $s_{0}$, equals to $\{s_{1},s_{2}\}$, but (3) $\mathsf{out}_{\mathtt{AG}}(s_{0},(\alpha_{1},\beta_{1}))$, the set of outcome states of $a$ and $b$ respectively performing $\alpha_{1}$ and $\beta_{1}$ at $s_{0}$, does not contain the intersection of $\mathsf{out}_{a}(s_{0},\alpha_{1})$ and $\mathsf{out}_{b}(s_{0},\beta_{1})$.
@@
  • $\mathtt{G}\textbf{-}\mathtt{AM}$ is independent if for all $s\in\mathtt{ST}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{{\mathrm{C}}}\in\mathtt{av}_{\mathrm{C}}(s)$, and $\sigma_{{\mathrm{D}}}\in\mathtt{av}_{\mathrm{D}}(s)$, $\sigma_{{\mathrm{C}}}\cup\sigma_{{\mathrm{D}}}\in\mathtt{av}_{\mathrm{C}\cup\mathrm{D}}(s)$;
@@
We let the three symbols $\mathtt{S}$, $\mathtt{I}$ and $\mathtt{D}$ correspond to the three properties, respectively, and let the eight strings $\epsilon$, $\mathtt{S}$, $\mathtt{I}$, $\mathtt{D}$, $\mathtt{SI}$, $\mathtt{SD}$, $\mathtt{ID}$, and $\mathtt{SID}$ correspond to the eight combinations of the three properties, respectively. We use $\mathrm{ES}$ to indicate the set of the eight strings.
For every $\mathtt{X}\in\mathrm{ES}$ and grand-coalition-first action model $\mathtt{G}\textbf{-}\mathtt{AM}$, we say $\mathtt{G}\textbf{-}\mathtt{AM}$ is a $\mathtt{X}$-model if $\mathtt{G}\textbf{-}\mathtt{AM}$ has the properties corresponding to $\mathtt{X}$.
For every $\mathtt{X}\in\mathrm{ES}$, we use $\mathsf{MCL+X}$ to refer to the coalition logic determined by the class of grand-coalition-first action $\mathtt{X}$-models.
@@ 3.2 Eight axiomatic systems
Independence of agents ($\mathtt{A}\text{-}\mathtt{IA}$): $(\langle[\mathrm{C}]\rangle\phi\land\langle[\mathrm{D}]\rangle\psi)\rightarrow\langle[\mathrm{C}\cup\mathrm{D}]\rangle(\phi\land\psi)$, where $\mathrm{C}\cap\mathrm{D}=\emptyset$
@@
For every $\mathtt{X}$ in $\mathrm{ES}$, the axiomatic system for $\mathsf{MCL+X}$ consists of the axioms and inference rules of $\mathsf{MCL}$, and the axioms corresponding to the elements of $\mathtt{X}$.
@@
For every $\mathtt{X}$ in $\mathrm{ES}$, the axiomatic system for $\mathsf{MCL+X}$ given in Definition 13 is sound and complete with respect to the set of valid formulas of $\mathsf{MCL+X}$.
@@
The class of grand-coalition-first $\mathsf{SID}$-models is $\alpha$-representable ([11, 5]), and the class of grand-coalition-first $\mathsf{SD}$-models is $\alpha$-representable ([14]). Are the other six classes of grand-coalition-first action models $\alpha$-representable? Are the eight classes of grand-coalition-first action models $z$-representable? All this is yet unknown.
@@ 4.1 Outcome functions are compositional in many situations
We say outcome functions of a class of action models are compositional if there is a way such that for every action model in the class, the outcome function of a coalition is determined by the outcome functions of its members in this way.
Outcome functions in grand-coalition-first action models are not compositional. Figure 2 gives a counterexample for this.
$s_{0}$$s_{1}$$s_{2}$$\{p,q\}$$\{p\}$$\{q\}$$\mathtt{G}\textbf{-}\mathtt{AM}_{1}$$(\alpha_{1},\beta_{1})\atop(\alpha_{2},\beta_{2})$$(\alpha_{1},\beta_{2})\atop(\alpha_{2},\beta_{1})$$(\alpha_{1},\beta_{1})$$(\alpha_{1},\beta_{1})$
$s_{0}$$s_{1}$$s_{2}$$\{p,q\}$$\{p\}$$\{q\}$$\mathtt{G}\textbf{-}\mathtt{AM}_{2}$$(\alpha_{1},\beta_{2})\atop(\alpha_{2},\beta_{1})$$(\alpha_{1},\beta_{1})\atop(\alpha_{2},\beta_{2})$$(\alpha_{1},\beta_{1})$$(\alpha_{1},\beta_{1})$
Figure 2: This figure offers a counterexample that outcome functions in grand-coalition-first action models are not generally compositional. We suppose $\mathtt{AG}=\{a,b\}$. Two grand-coalition-first action models $\mathtt{G}\textbf{-}\mathtt{AM}_{1}$ and $\mathtt{G}\textbf{-}\mathtt{AM}_{2}$ are depicted above, where $\{\mathsf{out}^{1}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ is the set of outcome functions of $\mathtt{G}\textbf{-}\mathtt{AM}_{1}$ and $\{\mathsf{out}^{2}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ is the set of outcome functions of $\mathtt{G}\textbf{-}\mathtt{AM}_{2}$. In $\mathtt{G}\textbf{-}\mathtt{AM}_{1}$: $\mathsf{out}^{1}_{a}(s_{0},\alpha_{1})=\{s_{1},s_{2}\},\mathsf{out}^{1}_{b}(s_{0},\beta_{1})=\{s_{1},s_{2}\}$, and $\mathsf{out}^{1}_{\mathtt{AG}}(s_{0},(\alpha_{1},\beta_{1}))=\{s_{1}\}$ In $\mathtt{G}\textbf{-}\mathtt{AM}_{2}$: $\mathsf{out}^{2}_{a}(s_{0},\alpha_{1})=\{s_{1},s_{2}\},\mathsf{out}^{2}_{b}(s_{0},\beta_{1})=\{s_{1},s_{2}\}$, and $\mathsf{out}^{2}_{\mathtt{AG}}(s_{0},(\alpha_{1},\beta_{1}))=\{s_{2}\}$. Note: $\mathsf{out}^{1}_{a}(s_{0},\alpha_{1})=\mathsf{out}^{2}_{a}(s_{0},\alpha_{1})$, $\mathsf{out}^{1}_{b}(s_{0},\beta_{1})=\mathsf{out}^{2}_{b}(s_{0},\beta_{1})$, but $\mathsf{out}^{1}_{\mathtt{AG}}(s_{0},(\alpha_{1},\beta_{1}))\neq\mathsf{out}^{2}_{\mathtt{AG}}(s_{0},(\alpha_{1},\beta_{1}))$. This implies that outcome functions in grand-coalition-first action models are not generally compositional: there is no way such that for every grand-coalition-first action model, the outcome function of a coalition is determined by the outcome functions of its members in this way.
There are many situations where every agent can change a component of the state, and different components of the state that can be changed by different agents are independent of each other. In these situations, the set of outcomes of a joint action is the intersection of the sets of outcomes of individual actions of the joint action.
@@
Adam and Bob control a ship lock with two doors: a front door and a back door. Adam can make the front door closed or open, Bob can make the back door closed or open, but they cannot make the two doors open at the same time. This situation can be represented by the grand-coalition-first action model depicted in Figure 3, where the set of outcomes of a joint action is the intersection of the sets of outcomes of individual actions of the joint action.
$s_{1}$$s_{2}$$s_{3}$$\{c_{f},c_{b}\}$$\{c_{f}\}$$\{c_{b}\}$$(\mathtt{skip},\mathtt{skip})$$(\mathtt{skip},\mathtt{skip})$$(\mathtt{skip},\mathtt{skip})$$(\mathtt{skip},\mathtt{open}\text{-}\mathtt{b})$$(\mathtt{skip},\mathtt{close}\text{-}\mathtt{b})$$(\mathtt{open}\text{-}\mathtt{f},\mathtt{skip})$$(\mathtt{close}\text{-}\mathtt{f},\mathtt{skip})$
Figure 3: This figure indicates a grand-coalition-first action model for the situation of Example 1. Here, $c_{f}$ and $c_{b}$, respectively, express the front door is closed and the back door is closed. It can be checked that in this model, the set of outcomes of a joint action is the intersection of the sets of outcomes of individual actions of the joint action. For example: (1) $\mathsf{out}_{\mathtt{AG}}(s_{1},(\mathtt{skip},\mathtt{skip}))=\mathsf{out}_{a}(s_{1},\mathtt{skip})\cap\mathsf{out}_{b}(s_{1},\mathtt{skip})=\{s_{1},s_{2}\}\cap\{s_{1},s_{3}\}=\{s_{1}\}$; (2) $\mathsf{out}_{\mathtt{AG}}(s_{2},(\mathtt{skip},\mathtt{close}\text{-}\mathtt{b}))=\mathsf{out}_{a}(s_{2},\mathtt{skip})\cap\mathsf{out}_{b}(s_{2},\mathtt{close}\text{-}\mathtt{f})=\{s_{1},s_{2}\}\cap\{s_{1}\}=\{s_{1}\}$; (3) $\mathsf{out}_{\{a,b\}}(s_{3},(\mathtt{close}\text{-}\mathtt{f},\mathtt{skip}))=\mathsf{out}_{a}(s_{3},\mathtt{close}\text{-}\mathtt{f})\cap\mathsf{out}_{b}(s_{3},\mathtt{skip})=\{s_{1}\}\cap\{s_{1},s_{3}\}=\{s_{1}\}$.
@@
A system has two processes, $a$ and $b$. The process $a$ can assign $1$ or 00 to the variable $x$, and the process $b$ can assign $1$ or 00 to the variable $y$. When $x=0$, $a$ can leave the value of $x$ unchanged or change it to $1$. When $x=1$, $a$ leaves the value of $x$ unchanged. Similarly, when $y=0$, $b$ can leave the value of $y$ unchanged or change it to $1$. When $y=1$, $b$ leaves the value of $y$ unchanged. Figure 4 indicates a grand-coalition-first model for this scenario, where the set of outcomes of a joint action is the intersection of the sets of outcomes of individual actions of the joint action.
@@
Figure 4: This figure indicates a grand-coalition-first action model for the situation of Example 2. It can be checked that in this model, the set of outcomes of a joint action is the intersection of the sets of outcomes of individual actions of the joint action. For example: (1) $\mathsf{out}_{\mathtt{AG}}(s_{1},(\mathtt{skip},\mathtt{skip}))=\mathsf{out}_{a}(s_{1},\mathtt{skip})\cap\mathsf{out}_{b}(s_{1},\mathtt{skip})=\{s_{1}\}\cap\{s_{1}\}=\{s_{1}\}$; (2) $\mathsf{out}_{\mathtt{AG}}(s_{2},(\mathtt{skip},\mathtt{y:=1}))=\mathsf{out}_{a}(s_{2},\mathtt{skip})\cap\mathsf{out}_{b}(s_{2},\mathtt{y:=1})=\{s_{1},s_{2}\}\cap\{s_{1}\}=\{s_{1}\}$; (3) $\mathsf{out}_{\mathtt{AG}}(s_{4},(\mathtt{x:=1},\mathtt{y:=1}))=\mathsf{out}_{a}(s_{4},\mathtt{x:=1})\cap\mathsf{out}_{b}(s_{4},\mathtt{y:=1})=\{s_{1},s_{2}\}\cap\{s_{1},s_{3}\}=\{s_{1}\}$.
We want to mention that some literature, such as [6] and [7], deals with agency by propositional control: there are some agents; every atomic proposition is controlled by at most one agent. In these settings, outcome functions are compositional.